The recommended pediatric IV dosage of Lasix is 1 mg/kg not to exceed 6 mg/kg. The child weighs 55 lbs. Lasix is available in 10 mg/ml. What is the volume to be administered IV initially?
Answer #1

recommended initial dose= 1mg/kg

recommended maximum dose= 6mg/kg

weight of the child= 55 lbs

1 lb= 0.453 kg

so, 55 lbs= 55* 0.453 kg= 24.915= 24.92 kg( weight of the child in kg)

available dose = 10 mg/ml

so initial dose = 1 mg/kg,

initial dose for 24.92 kg child =1* 24.92 kg = 24.92 mg

1ml = 10 mg

so 24.92 mg= X ml

(proportion method )

10 mg/1ml= 24.92 mg/X ml

10 X= 24.92

therefore, X= 24.92/ 10=2.492 ml

thus, 2.492 ml should be administered as the initial dose of lasix for 24.92 kg child.